Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-05-10
ISS On-Orbit Status 05/10/08

Saturday - day off for CDR Volkov, FE-1 Kononenko and FE-2 Reisman.

The crew performed the regular weekly three-hour task of thorough station cleaning. ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

As part of the house cleaning, Volkov & Kononenko performed regular maintenance cleaning on the fan grilles FS5, FS6, VPkhO, VdPrK, VPrK & TsV2 in the DC1 Docking Module and FGB (Funktsionalnyi-Grusovoi Blok).

The CDR conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

Sergey also gathered weekly data on Total Operating Time & 'On' durations of the Russian POTOK-150MK (150 micron) air filter unit of the SM's SOGS air revitalization subsystem for reporting to TsUP.

At ~10:00am EDT, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), RED resistive exerciser (FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

Later, Garrett downloaded the crewmembers' exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Working off the discretionary 'time permitting' task list, Volkov completed another EKON KPT-3 session, making observations and taking aerial photography of environmental conditions for Russia's Environmental Safety Agency (EKON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.

Also from the suggestions list, Kononenko performed a session of the Russian GFI-8 "Uragan" (hurricane) earth-imaging program, using the D2X to take telephotos of the western slope of the Pyrenees, the Chaiten volcano eruption in Chile as well as of the Kurils Islands in nadir.

A third job on the discretionary job list for the two cosmonauts, as per request from the European partner, still is to take one or two photographs of a Russian crewmember inside the ATV 'Jules Verne'.

Newly added to the 'available time' list was a request to take close-up photography of the two SKK materials exposure payloads mounted on the outside of the station,- SKK-2 on the 'Pirs' Docking Compartment (DC-1) from SM window #6, and SKK-9 on the SM hull, from the DC-1 EVA hatch window.

CDRA Update: The Carbon Dioxide Removal Assembly, running since 5/7, was deactivated yesterday by ground commanding after data from the ANITA (Analyzing Interferometer for Ambient Air) instrument showed that CDRA wasn't helping to remove the Freon-218 from the cabin air. The likely cause is that the pore size of the Zeolite bed in CDRA is smaller than Freon molecules, thus preventing the refrigerant from being adsorbed. (The Freon spill (~600 g) occurred on 4/29 after the replacement of the SKV-2 air conditioner compressor. To clean up quickly, the Russian BMP (Russian Harmful Impurities Removal System) was moded to regenerate its absorbent beds every 5 days instead of the regular 20 days. Freon-218 (Octafluoropropane/C3F8, Russian: Khladon) is safe (low toxicity, perhaps some irritation) and noncorrosive if not heated above ~600 degC. Primary hazard is oxygen displacement, as witness the Ozone Layer, but there is not enough C3F8 on board the ISS to significantly deplete any atmospheric oxygen.)

Water Dump Update: The US condensate water venting planned for 5/18 has been cancelled, as alternative options are being worked (such as storing the surplus water in clean ATV tanks).

PPS P6 Reconditioning Update: The reconditioning (discharge/recharge) maintenance of the P6 battery sets 2B3 and 4B3 was completed on 5/7, and both were returned to nominal operations. (P6 battery sets 2B2 & 4B2 were reconditioned last January.)

MSG Update: The annual re-certification of the Microgravity Science Glovebox was successfully completed.

Weekly Science Update (Expedition Seventeen -- Week 3)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Inc16 science data returned on 15S, waiting for hand-over to science team.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): B3/B4 plants are growing. A1/B1 and A2/B2 swapping was done, but the attempt to hydrate A1 and A2 in the B1 and B2 position is not currently successful.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed. Inc16 science data returned on 15S, waiting for hand-over to science team.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. - DEBIE-2: Link error still under investigation;- DOSTEL: On-going science acquisition;- EuTEMP: Currently inactive as planned; EVC: Further troubleshooting is on-going; Still problems to get high-rate data downlink; - EXPOSE: On-going science acquisition; - FIPEX: New science acquisition script has been restarted successfully on 4/30 and terminated nominally on 5/6. A new script was uplinked on 5/7 and is running since then; - MEDET: On-going science acquisition, running nominal; PLEGPAY: Instrument is powered on, but not in science acquisition mode; TRIBOLAB: On 4/25, the very sensitive TRIBOLAB Pin On Disk#2 (POD#2) run could be restarted and ran nominally until 5/9. The instrument has been put in Thermal Stabilization Mode (TSM) to prepare for an attempt to restart the Pin On Disk#1 (POD#1) run.

FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. 'Thank you Garrett for your great work! Teams continue to analyze on ground the Telemetry Logs, and further troubleshooting activities will be required with CEM-U (Upper) Optical Module.'

GEOFLOW: Start of GEOFLOW is pending further FSL troubleshooting and remaining commissioning activities (check of FSL optical modes).

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Garrett you have a couple of Sleep Activities scheduled for next week. You will have your monthly download and your last week of required Sleep logs scheduled as well. Any additional Sleep logging is above and beyond and greatly appreciated by the PI. After these activities are complete, you will have one more download and an Actiwatch doff on the Shuttle as your remaining activities.'

SOLAR (Solar Monitoring Observatory): Initial Sun tracking was started on 4/29, but the Sun was not in the proper Beta range to allow for nominal data acquisition, i.e. SOLAR platform went to Stand-By Mode when trying to track the Sun. On 4/30, solar measurements started with SOVIM and SOLSPEC instruments with regular acquisition of spectra and calibration runs. During this Sun observation period, the SOLAR platform, SOVIM and SOLSPEC Sun sensors positions have been correlated during a so-called 'criss-cross' procedure manoeuvre. The current Sun observation period will end during this weekend. - SOVIM: Acquiring science data - some glitches observed in the instrument telemetry; - SOLSPEC: Acquiring science data; - SOLACES: Acquiring science data, some recurrent synchronization problems between its microcontrollers.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. The first troubleshooting activity, which consists to insert back 6 Reference ECs into the Rotor A, is currently planned for next week.

CEO (Crew Earth Observations): Through 5/5 the ground has received a total of 982 of ISS/CEO images for review and cataloging. 'We are currently reviewing imagery with camera times corresponding to the following target list requests: Santa Barbara Coast, California - preliminarily acquired; and the Madrean Sky Islands, Mexico - not acquired. We have seen a number of photos suggesting your attempts to acquire the cities you requested, especially for Cologne, Germany - but there is no success so far to report. Your recent imagery of the New Orleans-Mississippi River-Lake Pontchartrain DYNAMIC EVENT (opening of the Bonnet Carre Spillway upriver from the city) will be published on NASA/GSFC's Earth Observatory website this weekend. Thanks for a great context view of this man-made event!'

CEO photo targets uplinked for today were Ouarkziz Impact Crater, Algeria (this 3.5-km in diameter impact structure stands out among the folded rocks of the Ouarkziz Monocline due to its well-defined circular crater. ISS had a nadir pass directly over the impact structure. Overlapping frames, using the 400 mm lens, acquired as ISS traversed the mountains from NW to SE were suggested to image both the crater and surrounding geologic structure), and Galeras Volcano, Columbia (the Galeras stratovolcano is located immediately west of the city of Pasto, and is one of Columbia's most active volcanoes. Numerous small to moderate eruptions have taken place since the first Spanish conquistadores arrived. ISS had a nadir pass directly over the volcano, and some cloud cover may have been present. Imagery of the volcano and surrounding deposits was requested).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use